OpenStudy (anonymous):

5d - d - 2d + 8 - 3d = 0 having trouble solving that

OpenStudy (uri):

Combine all d values.

OpenStudy (anonymous):

What do I do after combining like terms?

OpenStudy (uri):

Just combine them,Bring all d values together and add/subract them.

OpenStudy (anonymous):

(5-1-2-3)d + 8 =0 -d = -8 d=8

OpenStudy (uri):

@E.ali Try explaining them and let them find the answer..

OpenStudy (anonymous):

When I combined the terms I got: -1d + 8 = 0

OpenStudy (anonymous):

The answer is the explaining there !

OpenStudy (uri):

@yankeeez You're right Now subract 8 both sides -1d + 8 = 0 -8 -8 ============

OpenStudy (anonymous):

so... -1d = -8

OpenStudy (uri):

Right now divide -8 with -1 \[\frac{ -8 }{ -1 }\]

OpenStudy (uri):

double negative=positive.

OpenStudy (anonymous):

Yeah so d=8 now I got it thanks!

OpenStudy (uri):

@yankeeez yw :)